a shell vertically fired from earth with velocity v/n where v is escape velocity and n is some no. greater than 1neglecting the air resistance calculate the max. altitude attained by shell???

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Take this hint:

Apply Pythagoras’ theorem to the right-angled triangle to establish that the appropriate speed for a circular orbit just above the earth’s atmosphere is given by   

 

(Use the approximation that the distance traveled in one second is tiny compared to the radius of the earth.)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

vikas askiitian expert
13 years ago

in case of conservative forces we can apply conservation of energy ,so

initial total energy (TE)= PE + KE

PE = -GMm/R      (at surface)

KE=mv2/2n2      

v = (2GM/R)1/2   (escape speed)         so

KE = GMm/n2R

so total energy = KE + PE = GMm/n2R - GMm/R                  ..................1

finally  particle reaches to max height (x distance from surface)& its velocity becomes 0 ...

TE = PE + KE

PE = -GMm/(R+x) 

KE = 0

TE = -GMm/(R+x)           ....................2

since energy is conserved so eq 1 = eq 2

-GMm/(R+x) = -GMm/R + GMm/n2R

 X = R/(n2-1)                   (max altitude)

approve my ans if u like it